A chemist mixed a 15% glucose solution with a 35% glucose solution.This mixture produced 35 liters of a 19% glucose solution. How many liters of each solution did the chemist use in this mixture?

Answer by mananth(16946)

A chemist mixed a 15% glucose solution with a 35% glucose solution.This mixture produced 35 liters of a 19% glucose solution. How many liters of each solution did the chemist use in this mixture?
Let x be the quantity of 15% solution mixed
Let y be the quantity of 35% solution mixed
(15/100)x+(35/100)y = (19/100)35
0.15x+0.35y=0.19*35=6.65 ----1
x+y=35
Multiply this equation by 0.15
0.15x+0.15y=5.25 -----2
substract 2 from 1
0.2y=1.40
y= 7 liters
x=35-7=28
the quantity of 15% solution mixed = 28 liters
the be the quantity of 35% solution mixed = 7 liters
M ananth ------------ mananth@hotmail.com